OpenStudy (anonymous):

A manufacturing machine cuts a part at a rate of 150 feet per minute. What is the equivalent rate in centimeters per second?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

1 feet=30.48 centimeters 1 minute=60 seconds

OpenStudy (alexandervonhumboldt2):

convert: 150 feet => 150/30.48 centi 1 min => 1*60 sec

OpenStudy (alexandervonhumboldt2):

apromaxietly 4.9 centi per 60 seconds

OpenStudy (anonymous):

so do I multiply them:)?

OpenStudy (alexandervonhumboldt2):

well you have to multiply 1 min by 60 sec to get conert of 1 min=60sec. when you convert feet to centi you have to divide as centi is < than feet.

OpenStudy (anonymous):

ok. so can you give me the formula to use as an example

OpenStudy (alexandervonhumboldt2):

so you want a formula as an example to convert some unit to another?

OpenStudy (alexandervonhumboldt2):

so the formula for an example to convert feets to centi is f=30.48centi

OpenStudy (anonymous):

no, the formula for this one:)

OpenStudy (alexandervonhumboldt2):

hmm there is no exact formula for converting all units. each has it's own.

OpenStudy (anonymous):

I got an answer of 76.2 cm per second in the answer book. Im just trying to figure out how they got that answer.

OpenStudy (alexandervonhumboldt2):

oh sorry theey ask per 1 second not per 60 seeconds. let me redo.

OpenStudy (anonymous):

okay:)

OpenStudy (alexandervonhumboldt2):

150 feet => 150*30.48 centi 1 munite=60 seocnds. first: divide 150*30.48.=4572 now divide this by by 60: 4572/60=76.2 here is the correct answer

OpenStudy (alexandervonhumboldt2):

first multiply i meant

OpenStudy (anonymous):

awesome!!!